# PsychoBench: Evaluating the Psychology Intelligence of Large Language Models

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable success across a wide range of industries, primarily due to their impressive generative abilities. Yet, their potential in applications requiring cognitive abilities, such as psychological counseling, remains largely untapped. This paper investigates the key question: Can LLMs be effectively applied to psychological counseling? To determine whether an LLM can effectively take on the role of a psychological counselor, the first step is to assess whether it meets the qualifications required for such a role, namely the ability to pass the U.S. National Counselor Certification Exam (NCE). This is because, just as a human counselor must pass a certification exam to practice, an LLM must demonstrate sufficient psychological knowledge to meet the standards required for such a role. To address this, we introduce PsychoBench, a benchmark grounded in U.S.national counselor examinations, a licensure test for professional counselors that requires about 70% accuracy to pass. PsychoBench comprises approximately 2,252 carefully curated single-choice questions, crafted to require deep understanding and broad enough to cover various sub-disciplines of psychology. This benchmark provides a comprehensive assessment of an LLM’s ability to function as a counselor. Our evaluation shows that advanced models such as GPT-4o, Llama3.3-70B, and Gemma3-27B achieve well above the passing threshold, while smaller open-source models (e.g., Qwen2.5-7B, Mistral-7B) remain far below it. These results suggest that only frontier LLMs are currently capable of meeting counseling exam standards, highlighting both the promise and the challenges of developing psychology-oriented LLMs. We release the proposed dataset for public use: https://github.com/cloversjtu/PsychoBench
PsychoBench: Evaluating the Psychology Intelligence of Large Language Models
@@ 1 Introduction
Psychological counseling is a profession that requires comprehensive knowledge of mental health theories, counseling techniques, ethics, and case analysis skills Zhang and Wang (2019); Chen and Liu (2019). Licensed counselors must pass a national certification examination to demonstrate competence across these areas Wang and Zhao (2023), which typically includes carefully designed questions spanning multiple subfields of psychology and counseling practice Wang and Zhao (2020).
@@
With the rapid advancement of LLMs, their potential application in specialized professional domains, such as psychological counseling, has become increasingly significant. A key question is whether an LLM can effectively perform the tasks of a licensed psychological counselor. A natural first step toward answering this question is to evaluate whether an LLM can pass the national certification exam for psychological counselors. However, there is currently a lack of publicly available exam questions to measure LLMs’ ability to succeed in this evaluation. Although LLMs have demonstrated remarkable progress in natural language understanding and generation, excelling in tasks such as question answering, dialogue systems, and reasoning Bai et al. (2023), most evaluations have focused on general knowledge, mathematical reasoning, and coding tasks Chen et al. (2021); Li and Zhao (2022). Assessing LLMs in the context of psychological counseling is therefore essential to determine their potential to assist human counselors and support educational training Smith and Lee (2023); Zhao and Chen (2024). Addressing this gap is important not only for AI research but also for education, mental health support, and social science studies.
@@
To address this need, we introduce PsychoBench, a benchmark designed to evaluate LLMs’ capability to perform psychological counseling tasks. PsychoBench comprises approximately 2,252 single-choice questions that were initially collected from publicly available psychological counseling exam questions on the internet. To enhance linguistic diversity and clarity, the questions were paraphrased and refined using GPT-based methods Xu et al. (2024b). Each question was then carefully reviewed and corrected by human experts to ensure accuracy and consistency, providing a high-quality benchmark covering multiple subfields, including counseling methods, abnormal psychology, developmental psychology, and ethical considerations Wang and Zhao (2020); Chen and Liu (2019).
@@
Our evaluation of state-of-the-art LLMs on PsychoBench shows that these models achieve relatively high accuracy, demonstrating promising potential to understand and reason about psychological knowledge and counseling scenarios OpenAI (2023). Nevertheless, a notable performance gap remains compared to expert human professionals, particularly in nuanced understanding, contextual judgment, and ethical reasoning. This underscores the need for further research to enhance LLMs’ competencies in emotionally sensitive, ethically grounded, and personalized counseling tasks.
@@ 2.1 LLMs in Mental Health Care
Recent studies have explored the potential of Large Language Models (LLMs) in mental health tasks, including understanding mental health concepts, ethical reasoning, and counseling-related scenarios Smith and Lee (2023); Zhao and Chen (2024). While LLMs are capable of providing preliminary analysis and suggestions, they continue to fall short of human experts, particularly in areas requiring nuanced understanding, emotional reasoning, and contextual judgment.
Several systematic and scoping reviews have examined the application of LLMs in mental health care. For instance, Xu et al. (2024a) and Wang and Li (2024) have highlighted the promise of LLMs for generative tasks such as diagnosis and therapy, though they also acknowledge ongoing challenges related to emotional intelligence and personalized care. Chen and Liu (2024) assessed the capabilities of generative AI in mental health, noting its strengths in early diagnosis and data generation but also identifying significant gaps in emotional sensitivity and ethical reasoning. A review by Liu et al. (2023) specifically focused on BERT-based models for suicide detection and risk assessment, emphasizing the need for further improvements in accuracy for clinical applications. Additionally, Zhao and Chen (2024) reviewed the application of LLMs in psychological counseling, highlighting both their potential and the limitations that currently hinder their broader use.
@@ 2.2 LLMs in Psychological and Social Domains
Recent works have also highlighted the role of LLMs in psychological assessment and therapy. Brown and Peterson (2023) examined the use of LLMs for virtual therapy, showcasing their potential to assist therapists in providing empathetic responses and therapeutic interventions. Liu and Zhang (2023) explored how LLMs can be used to detect emotions in text-based dialogues, improving the assessment of mental health conditions like depression or anxiety.
In the realm of crisis intervention, Johnson and Miller (2023) reviewed LLM applications in suicide prevention, demonstrating how AI models can recognize early signs of self-harm in text-based communication. Additionally, Miller and Davis (2024) explored how AI systems can be used for real-time crisis intervention, aiding professionals in making timely decisions during emergencies.
Ethical considerations remain a significant concern in the use of LLMs for mental health tasks. Jones and Thompson (2023) discussed the ethical implications, focusing on privacy, informed consent, and the risk of AI-generated misinformation. Ensuring that LLMs adhere to ethical standards is vital for their integration into professional mental health practices Smith and Johnson (2023).
Finally, the collaborative potential between LLMs and human counselors has been explored as a means to enhance therapeutic outcomes. Taylor and Green (2024) highlighted how human-AI collaboration could lead to more personalized care, where LLMs assist human professionals by offering additional insights and suggestions. These collaborations aim to create a more holistic approach to mental health support. Other psychology-related datasets have also been introduced in adjacent domains, such as PsyMo (Cosma and Radoi, 2024), which focuses on estimating psychological traits from gait data. However, these datasets target computer vision tasks and are orthogonal to our goal of evaluating LLMs’ psychological counseling competence.
3 PsychoBench Dataset
@@
In total, approximately 2,252 questions were included in the final dataset. All questions are anonymized and stripped of any identifying information. The dataset is made publicly available on GitHub [^1] https://github.com/cloversjtu/PsychoBench under a CC-BY-NC-ND license.
@@ 3.1 Collecting Data
PsychoBench Example Item Question: Juanita is well-liked by her peers due to her sociable, relaxed, and energetic nature. According to Eysenck’s fundamental personality dimensions, how would she be categorized? Options: 1. Extroverted and stable  (Correct) 2. Passive-aggressive 3. Intrinsically motivated 4. Introverted and unstable 5. Cyclothymic and dysthymic Rationale: According to Eysenck’s two fundamental personality dimensions, sociable and energetic behaviors reflect high extraversion, while being relaxed and well-liked indicates emotional stability. Therefore, Juanita is best categorized as extroverted and stable.
@@
Figure 1: Overview of the PsychoBench task, which evaluates LLMs on counselor exam-style multiple-choice items (left) using the pipeline illustrated on the right.
@@ 3.2 Dataset Statistics
We provide a detailed overview of PsychoBench to highlight its scale, diversity, and quality. Specifically, we analyze the distribution of questions across psychological subfields, the variation in question length and option count, and the balance between expert-authored and GPT-refined items. The final dataset contains a total of 2,252 multiple-choice questions. On average, each question consists of 23.62 words in the stem (measured by whitespace tokenization), reflecting a moderate level of linguistic complexity. In terms of option distribution, most questions provide either four or five candidate answers: 1,142 questions (50.71%) have four options, 1,102 questions (48.93%) have five options, while only 8 questions (0.36%) contain two options. Each question has a single correct answer, ensuring consistency with the format of standard psychological counseling certification exams. The detailed statistics are shown in Table 1.
@@ 3.3 Models Evaluated
Table 2: Open-source and closed-source models evaluated on PsychoBench.
@@ 3.4 Task Definition
PsychoBench is formulated as a multiple-choice question answering task. Each instance consists of a question stem and a set of candidate options, with exactly one option marked as correct. The number of candidate options varies across questions: most items provide either four or five options, while a small portion contains only two. Given a question $q$ and its associated options $\{o_{1},o_{2},...,o_{n}\}$, where $n\in\{2,4,5\}$, the task is to predict the correct answer $o^{*}\in\{o_{1},o_{2},...,o_{n}\}$. This formulation is inspired by the format of the U.S. National Counselor Examination (NCE), making it a natural proxy for assessing an LLM’s ability to acquire and apply psychological knowledge. We frame the task as a classification problem, where models must select the correct option among candidates. Evaluation is primarily based on accuracy, i.e., the proportion of correctly predicted items over the total number of questions.
@@
Table 3: Performance of different LLMs on the PsychoBench dataset.
@@ 4.1 Experimental Setup
We evaluate a diverse set of LLMs on the constructed psychological counseling benchmark dataset PsychoBench. The dataset contains 2,252 multiple-choice questions, each with 2–5 candidate options and exactly one correct answer. Models are required to predict the correct option given the question and its candidate choices.
@@ 4.2 Results
Table 3 presents the performance of a wide range of LLMs on the PsychoBench dataset. The results show that proprietary frontier models clearly lead the benchmark. GPT-4o achieves the highest Top-1 accuracy of 94.36%, with GPT-4o-mini (90.49%) and Llama3.3-70B-Instruct (91.16%) following closely. These models also maintain weighted F1, precision, and recall all above 90%, confirming their robustness across different question types. Such scores are well above the passing threshold of the U.S. National Counselor Examination (NCE), which requires roughly 70% accuracy, suggesting that cutting-edge proprietary systems already possess sufficient knowledge to meet licensure-level requirements.
@@ 5 Conclusions
This work introduces PsychoBench, the first benchmark designed to systematically evaluate the psychological counseling competence of large language models. By grounding the evaluation in the U.S. National Counselor Examination (NCE), the benchmark provides a high-stakes, practice-relevant standard for assessing whether LLMs possess the knowledge required for professional counseling. Our experimental results reveal a clear performance hierarchy: cutting-edge proprietary models such as GPT-4o and GPT-4o-mini already exceed licensure-level requirements, large open-source systems like Llama3.3-70B and Gemma3-27B approach this level but still lag slightly behind, while mid-sized and small-scale models remain unreliable. These findings highlight both the promise and limitations of current LLMs in psychology-oriented applications.
Overall, PsychoBench underscores that while frontier LLMs demonstrate strong mastery of exam-style psychological knowledge, substantial work remains before smaller open-source systems can achieve reliable competency. We view this benchmark not only as an evaluation tool but also as a call to action: advancing specialized training, domain adaptation, and reasoning calibration is essential for developing safe, trustworthy, and accessible AI systems in mental health contexts.
@@ 6 Limitations
While PsychoBench provides the first systematic benchmark for evaluating psychological counseling competence in large language models, several limitations remain. First, the benchmark is constructed from multiple-choice questions in the U.S. National Counselor Examination (NCE), which primarily assess factual and applied psychological knowledge but may not fully capture the interpersonal, empathic, and conversational skills required in real-world counseling. Passing the exam is a necessary but not sufficient condition for competent counseling practice. Second, our benchmark is limited to English and U.S.-specific licensure standards, which may restrict its generalizability across cultural and linguistic contexts. Third, our evaluation focuses on answer selection accuracy and does not assess other important aspects such as reasoning transparency, calibration of confidence, or the ability to handle ambiguous or emotionally charged scenarios. Finally, although we evaluate a diverse range of proprietary and open-source models, the rapid evolution of LLMs means that newer architectures may quickly change the performance landscape.

### v4

# PsychCounsel-Bench: Evaluating the Psychology Intelligence of Large Language Models

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable success across a wide range of industries, primarily due to their impressive generative abilities. Yet, their potential in applications requiring cognitive abilities, such as psychological counseling, remains largely untapped. This paper investigates the key question: Can LLMs be effectively applied to psychological counseling? To determine whether an LLM can effectively take on the role of a psychological counselor, the first step is to assess whether it meets the qualifications required for such a role, namely the ability to pass the U.S. National Counselor Certification Exam (NCE). This is because, just as a human counselor must pass a certification exam to practice, an LLM must demonstrate sufficient psychological knowledge to meet the standards required for such a role. To address this, we introduce PsychCounsel-Bench, a benchmark grounded in U.S.national counselor examinations, a licensure test for professional counselors that requires about 70% accuracy to pass. PsychCounsel-Bench comprises approximately 2,252 carefully curated single-choice questions, crafted to require deep understanding and broad enough to cover various sub-disciplines of psychology. This benchmark provides a comprehensive assessment of an LLM’s ability to function as a counselor. Our evaluation shows that advanced models such as GPT-4o, Llama3.3-70B, and Gemma3-27B achieve well above the passing threshold, while smaller open-source models (e.g., Qwen2.5-7B, Mistral-7B) remain far below it. These results suggest that only frontier LLMs are currently capable of meeting counseling exam standards, highlighting both the promise and the challenges of developing psychology-oriented LLMs. We release the proposed dataset for public use: https://github.com/cloversjtu/PsychCounsel-Bench
PsychCounsel-Bench: Evaluating the Psychology Intelligence of Large Language Models
@@ 1 Introduction
Psychological counseling is a profession that requires comprehensive knowledge of mental health theories, counseling techniques, ethics, and case analysis skills zhang2019psychological; chen2019ethics. Licensed counselors must pass a national certification examination to demonstrate competence across these areas wang2023psychologyexam, which typically includes carefully designed questions spanning multiple subfields of psychology and counseling practice wang2020psychosubfields.
@@
With the rapid advancement of LLMs zeng2019dirichlet; ji2023rho; zeng2024dirichlet; zeng2025task, their potential application in specialized professional domains, such as psychological counseling, has become increasingly significant. A key question is whether an LLM can effectively perform the tasks of a licensed psychological counselor. A natural first step toward answering this question is to evaluate whether an LLM can pass the national certification exam for psychological counselors. However, there is currently a lack of publicly available exam questions to measure LLMs’ ability to succeed in this evaluation. Although LLMs have demonstrated remarkable progress in natural language understanding and generation, excelling in tasks such as question answering, dialogue systems, and reasoning bai2023scaling, most evaluations have focused on general knowledge, mathematical reasoning, and coding tasks chen2021evaluating; li2022mathematical. Assessing LLMs in the context of psychological counseling is therefore essential to determine their potential to assist human counselors and support educational training smith2023psychological; zhao2024psychological. Addressing this gap is important not only for AI research but also for education, mental health support, and social science studies.
@@
To address this need, we introduce PsychCounsel-Bench, a benchmark designed to evaluate LLMs’ capability to perform psychological counseling tasks. PsychCounsel-Bench comprises approximately 2,252 single-choice questions that were initially collected from publicly available psychological counseling exam questions on the internet. To enhance linguistic diversity and clarity, the questions were paraphrased and refined using GPT-based methods xu2024datarefinement. Each question was then carefully reviewed and corrected by human experts to ensure accuracy and consistency, providing a high-quality benchmark covering multiple subfields, including counseling methods, abnormal psychology, developmental psychology, and ethical considerations wang2020psychosubfields; chen2019ethics.
@@
Our evaluation of state-of-the-art LLMs on PsychCounsel-Bench shows that these models achieve relatively high accuracy, demonstrating promising potential to understand and reason about psychological knowledge and counseling scenarios openai2023gpt4. Nevertheless, a notable performance gap remains compared to expert human professionals, particularly in nuanced understanding, contextual judgment, and ethical reasoning. This underscores the need for further research to enhance LLMs’ competencies in emotionally sensitive, ethically grounded, and personalized counseling tasks.
@@ 2.1 LLMs in Mental Health Care
Recent studies have explored the potential of Large Language Models (LLMs) in mental health tasks, including understanding mental health concepts, ethical reasoning, and counseling-related scenarios smith2023psychological; zhao2024psychological. While LLMs are capable of providing preliminary analysis and suggestions, they continue to fall short of human experts, particularly in areas requiring nuanced understanding, emotional reasoning, and contextual judgment.
Several systematic and scoping reviews have examined the application of LLMs in mental health care. For instance, xu2024survey and wang2024mentalhealthsurvey have highlighted the promise of LLMs for generative tasks such as diagnosis and therapy, though they also acknowledge ongoing challenges related to emotional intelligence and personalized care. chen2024counselingai assessed the capabilities of generative AI in mental health, noting its strengths in early diagnosis and data generation but also identifying significant gaps in emotional sensitivity and ethical reasoning. A review by liu2023mentalhealth specifically focused on BERT-based models for suicide detection and risk assessment, emphasizing the need for further improvements in accuracy for clinical applications. Additionally, zhao2024psychological reviewed the application of LLMs in psychological counseling, highlighting both their potential and the limitations that currently hinder their broader use.
@@ 2.2 LLMs in Psychological and Social Domains
Recent works have also highlighted the role of LLMs in psychological assessment and therapy. brown2023therapyai examined the use of LLMs for virtual therapy, showcasing their potential to assist therapists in providing empathetic responses and therapeutic interventions. liu2023emotionrecognition explored how LLMs can be used to detect emotions in text-based dialogues, improving the assessment of mental health conditions like depression or anxiety.
In the realm of crisis intervention, johnson2023suicideprevention reviewed LLM applications in suicide prevention, demonstrating how AI models can recognize early signs of self-harm in text-based communication. Additionally, miller2024crisisai explored how AI systems can be used for real-time crisis intervention, aiding professionals in making timely decisions during emergencies.
Ethical considerations remain a significant concern in the use of LLMs for mental health tasks. jones2023ethics discussed the ethical implications, focusing on privacy, informed consent, and the risk of AI-generated misinformation. Ensuring that LLMs adhere to ethical standards is vital for their integration into professional mental health practices smith2023ethicalAI.
Finally, the collaborative potential between LLMs and human counselors has been explored as a means to enhance therapeutic outcomes. taylor2024humanai highlighted how human-AI collaboration could lead to more personalized care, where LLMs assist human professionals by offering additional insights and suggestions. These collaborations aim to create a more holistic approach to mental health support. Other psychology-related datasets have also been introduced in adjacent domains, such as PsyMo (cosma2024psymo), which focuses on estimating psychological traits from gait data. However, these datasets target computer vision tasks and are orthogonal to our goal of evaluating LLMs’ psychological counseling competence.
3 PsychCounsel-Bench Dataset
@@
In total, approximately 2,252 questions were included in the final dataset. All questions are anonymized and stripped of any identifying information. The dataset is made publicly available on GitHub [^1] https://github.com/cloversjtu/PsychCounsel-Bench under a CC-BY-NC-ND license.
@@ 3.1 Collecting Data
PsychCounsel-Bench Example Item Question: Juanita is well-liked by her peers due to her sociable, relaxed, and energetic nature. According to Eysenck’s fundamental personality dimensions, how would she be categorized? Options: 1. Extroverted and stable  (Correct) 2. Passive-aggressive 3. Intrinsically motivated 4. Introverted and unstable 5. Cyclothymic and dysthymic Rationale: According to Eysenck’s two fundamental personality dimensions, sociable and energetic behaviors reflect high extraversion, while being relaxed and well-liked indicates emotional stability. Therefore, Juanita is best categorized as extroverted and stable.
@@
Figure 1: Overview of the PsychCounsel-Bench task, which evaluates LLMs on counselor exam-style multiple-choice items (left) using the pipeline illustrated on the right.
@@ 3.2 Dataset Statistics
We provide a detailed overview of PsychCounsel-Bench to highlight its scale, diversity, and quality. Specifically, we analyze the distribution of questions across psychological subfields, the variation in question length and option count, and the balance between expert-authored and GPT-refined items. The final dataset contains a total of 2,252 multiple-choice questions. On average, each question consists of 23.62 words in the stem (measured by whitespace tokenization), reflecting a moderate level of linguistic complexity. In terms of option distribution, most questions provide either four or five candidate answers: 1,142 questions (50.71%) have four options, 1,102 questions (48.93%) have five options, while only 8 questions (0.36%) contain two options. Each question has a single correct answer, ensuring consistency with the format of standard psychological counseling certification exams. The detailed statistics are shown in Table 1.
@@ 3.3 Models Evaluated
Table 2: Open-source and closed-source models evaluated on PsychCounsel-Bench.
@@ 3.4 Task Definition
PsychCounsel-Bench is formulated as a multiple-choice question answering task. Each instance consists of a question stem and a set of candidate options, with exactly one option marked as correct. The number of candidate options varies across questions: most items provide either four or five options, while a small portion contains only two. Given a question $q$ and its associated options $\{o_{1},o_{2},...,o_{n}\}$, where $n\in\{2,4,5\}$, the task is to predict the correct answer $o^{*}\in\{o_{1},o_{2},...,o_{n}\}$. This formulation is inspired by the format of the U.S. National Counselor Examination (NCE), making it a natural proxy for assessing an LLM’s ability to acquire and apply psychological knowledge. We frame the task as a classification problem, where models must select the correct option among candidates. Evaluation is primarily based on accuracy, i.e., the proportion of correctly predicted items over the total number of questions.
@@
Table 3: Performance of different LLMs on the PsychCounsel-Bench dataset.
@@ 4.1 Experimental Setup
We evaluate a diverse set of LLMs on the constructed psychological counseling benchmark dataset PsychCounsel-Bench. The dataset contains 2,252 multiple-choice questions, each with 2–5 candidate options and exactly one correct answer. Models are required to predict the correct option given the question and its candidate choices.
@@ 4.2 Results
Table 3 presents the performance of a wide range of LLMs on the PsychCounsel-Bench dataset. The results show that proprietary frontier models clearly lead the benchmark. GPT-4o achieves the highest Top-1 accuracy of 94.36%, with GPT-4o-mini (90.49%) and Llama3.3-70B-Instruct (91.16%) following closely. These models also maintain weighted F1, precision, and recall all above 90%, confirming their robustness across different question types. Such scores are well above the passing threshold of the U.S. National Counselor Examination (NCE), which requires roughly 70% accuracy, suggesting that cutting-edge proprietary systems already possess sufficient knowledge to meet licensure-level requirements.
@@ 5 Conclusions
This work introduces PsychCounsel-Bench, the first benchmark designed to systematically evaluate the psychological counseling competence of large language models. By grounding the evaluation in the U.S. National Counselor Examination (NCE), the benchmark provides a high-stakes, practice-relevant standard for assessing whether LLMs possess the knowledge required for professional counseling. Our experimental results reveal a clear performance hierarchy: cutting-edge proprietary models such as GPT-4o and GPT-4o-mini already exceed licensure-level requirements, large open-source systems like Llama3.3-70B and Gemma3-27B approach this level but still lag slightly behind, while mid-sized and small-scale models remain unreliable. These findings highlight both the promise and limitations of current LLMs in psychology-oriented applications.
Overall, PsychCounsel-Bench underscores that while frontier LLMs demonstrate strong mastery of exam-style psychological knowledge, substantial work remains before smaller open-source systems can achieve reliable competency. We view this benchmark not only as an evaluation tool but also as a call to action: advancing specialized training, domain adaptation, and reasoning calibration is essential for developing safe, trustworthy, and accessible AI systems in mental health contexts.
@@ 6 Limitations
While PsychCounsel-Bench provides the first systematic benchmark for evaluating psychological counseling competence in large language models, several limitations remain. First, the benchmark is constructed from multiple-choice questions in the U.S. National Counselor Examination (NCE), which primarily assess factual and applied psychological knowledge but may not fully capture the interpersonal, empathic, and conversational skills required in real-world counseling. Passing the exam is a necessary but not sufficient condition for competent counseling practice. Second, our benchmark is limited to English and U.S.-specific licensure standards, which may restrict its generalizability across cultural and linguistic contexts. Third, our evaluation focuses on answer selection accuracy and does not assess other important aspects such as reasoning transparency, calibration of confidence, or the ability to handle ambiguous or emotionally charged scenarios. Finally, although we evaluate a diverse range of proprietary and open-source models, the rapid evolution of LLMs means that newer architectures may quickly change the performance landscape.